A card is drawn from a standard deck of 52 cards. Find the probability that the card is either a six or a black card.
Answer by stanbon(57984)

A card is drawn from a standard deck of 52 cards. Find the probability that the card is either a six or a black card.
----
# of 6 or black cards = 4 + 26 - 2 = 28
----
P(6 or black) = 28/52 = 7/13
=================================
Cheers,
Stan H.
============